hiiiihiiiiiiiiii:

Add the fractions 3/8+1/8. Remember to reduce. 4/8 None of the answers is correct. 1/2 5/8

justus:

The denominator is the same so just add the numerators.

hiiiihiiiiiiiiii:

none

hiiiihiiiiiiiiii:

i thank it none

justus:

That's incorrect.

jimthompson5910:

Let's say we had a cake we cut into 8 slices 3/8 means 3 slices out of 8 1/8 means 1 slice out of 8 Adding 3/8+1/8 means we started with 3 slices and add on 1 slice 

hiiiihiiiiiiiiii:

its none of them right

jimthompson5910:



xxDeppressionxx:

Answer: 38+18=12 Solution with Steps 38+18=? The two fractions have like denominators so you can add the numerators. Then: 3+18=48 This fraction can be reduced by dividing both the numerator and denominator by the Greatest Common Factor of 4 and 8 using GCF(4,8) = 4 4÷48÷4=12 Therefore: 38+18=12 Solution by Formulas Apply the fractions formula for addition, to 38+18 and solve (3×8)+(1×8)8×8 =24+864 =3264 Reduce by dividing both the numerator and denominator by the Greatest Common Factor GCF(32,64) = 32 32÷3264÷32=12 Therefore: 38+18=12

xxDeppressionxx:

It is 3/8+1/8=1/2